OpenStudy (anonymous):

2. Given the function f(x) = log2(x + 6), find the value of f−1(3).

OpenStudy (er.mohd.amir):

what do u mean by f-1(3) ? may be something is nor right.

OpenStudy (anonymous):

f^-1 (3)

OpenStudy (anonymous):

sorry about that

OpenStudy (campbell_st):

you need to find the inverse function 1st before anything else... swap x and y and you have \[x = \log_{2}(y + 6)\] now make y the subject... once you have the inverse, just substitute x = 3 hope it helps

OpenStudy (anonymous):

Hi, I have no idea how to do that. do you have the time to help me?

OpenStudy (anonymous):

Here are the answers and the original question if that helps.

OpenStudy (campbell_st):

well there is a log law that says \[a^{\log_{a}(b)} = b\] which requires you to raise each side of the equation as a power of the base of 2 so \[2^x = 2^{\log_{2}(y + 6)}\] so what do you think happens next

OpenStudy (anonymous):

Would you try to get rid of the y+6?

OpenStudy (campbell_st):

well that is what is left, the fact that the base of the exponent and base of the log are the same, the problem becomes \[2^x = y + 6\] now solve for y. that will be the inverse function. last part to find \[f^{-1}(3)\] just substitute x = 3 into the function and evlauate

OpenStudy (anonymous):

so 8=y+6. And then what do you do for the f-1(3) part?

OpenStudy (campbell_st):

you are being asked to find the value of y that makes the equation true

OpenStudy (campbell_st):

\[y = 2^x - 6~~~or~~~~~f^{-1}(x) = 2^x - 6\]

OpenStudy (anonymous):

Ok. none of those are my answer choices though, in the pitcture i posted earlier are the answer choices. Sorry about this, I just do not get this problem at all

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well that is the equation of the inverse function.... you are asked to find the value of the equation when you substitute x = 3.

OpenStudy (campbell_st):

so make the substitution and get the answer.

OpenStudy (anonymous):

so f-1 (3)=2?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

that's correct